tickets at a concert cost $8 for section A and $4.25 for section B. In total 4500 tickets were sold worth $30000. How many of each type of ticket were sold?
Answer by jim_thompson5910(35256)

tickets at a concert cost $8 for section A and $4.25 for section B. In total 4500 tickets were sold worth $30000. How many of each type of ticket were sold?

-------------------------------------------------------

Let x = # of section A tickets sold
Let y = # of section B tickets sold

4500 tickets were sold, so x+y = 4500. Solve for y to get y = 4500-x

"tickets at a concert cost $8 for section A and $4.25 for section B. In total 4500 tickets were sold worth $30000" all translates to the equation below

8x + 4.25y = 30000

8x + 4.25(4500-x) = 30000 ... plug in y = 4500 - x

8x + 4.25(4500) + 4.25(-x) = 30000

8x + 19125 - 4.25x = 30000

3.75x + 19125 = 30000

3.75x = 30000 - 19125

3.75x = 10875

x = 10875/3.75

x = 2900

y = 4500 - x

y = 4500 - 2900

y = 1600

Since x = 2900 and y = 1600 this means that

2900 section A tickets were sold
1600 section B tickets were sold